OpenStudy (anonymous):

HOW THE HELL DO I ANSWER THIS?!?!?!?! 2x-7+3x=4x+2

OpenStudy (ledah):

well lets look at this. Simplifying 2x + 7 + 3x = 4x + -2 Reorder the terms: 7 + 2x + 3x = 4x + -2 Combine like terms: 2x + 3x = 5x 7 + 5x = 4x + -2 Reorder the terms: 7 + 5x = -2 + 4x Solving 7 + 5x = -2 + 4x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-4x' to each side of the equation. 7 + 5x + -4x = -2 + 4x + -4x Combine like terms: 5x + -4x = 1x 7 + 1x = -2 + 4x + -4x Combine like terms: 4x + -4x = 0 7 + 1x = -2 + 0 7 + 1x = -2 Add '-7' to each side of the equation. 7 + -7 + 1x = -2 + -7 Combine like terms: 7 + -7 = 0 0 + 1x = -2 + -7 1x = -2 + -7 Combine like terms: -2 + -7 = -9 1x = -9 Divide each side by '1'. x = -9 Simplifying x = -9

OpenStudy (ledah):

you could find these solutions with many steps. you should always aim to get x by itself. okay? hope this helps